OpenStudy (anonymous):

What is the surface area of the rectangular prism? 528 cm 2 440 cm 2 404 cm 2 484 cm 2

OpenStudy (michele_laino):

where is your prism, please?

OpenStudy (michele_laino):

hint: the requested total surface S is: \[\Large S = 2 \times \left( {11 \times 8 + 11 \times 6 + 6 \times 8} \right) = ...\]

OpenStudy (anonymous):

404

OpenStudy (michele_laino):

since we have two 8x11 rectangles, two 6x11 rectangles, and two 6x8 rectangles

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

xD